109.6 million is 1054% of what number?
Answer by nerdybill(7384)

109.6 million is 1054% of what number?
.
Let x = the number
then
109.6 million = 1054% * x
109.6 * 10^6 = 1054/100 * x
109.6 * 1000000 = 10.54 * x
109600000 = 10.54x
109600000/10.54 = x
10398481.973 = x
OR
10.398 million = x